OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. E(-2, 2), F(5, 1)

OpenStudy (anonymous):

find the slope over 7, down 1, slope is \(-\frac{1}{7}\) or use \[m=\frac{2-1}{-2-5}=-\frac{1}{7}\] then use point slope formula

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] with \[m=-\frac{1}{7},x_1=5,y_1=1\]

OpenStudy (anonymous):

ok, but how do I continue after I have \[y-1=\left(\begin{matrix}-1 \\7\end{matrix}\right)x+\left(\begin{matrix}5 \\7\end{matrix}\right)\] @satellite73 ?

OpenStudy (anonymous):

you do whats in parentheses the do whats left on the outside